Write the equation in standard form of the line that is perpendicular to the line y=-2x+6 and passes through the point (-4,8)
Answer by mananth(16946)

1 y = -2 x + 6
Divide by 1
y = -2 x + 6
Compare this equation with y=mx+b, m= slope & b= y intercept
slope m = -2

The slope of a line perpendicular to the above line will be the negative reciprocal 0.5
m1*m2=-1
The slope of the required line will be 1/2

m= 1/2 ,point ( -4 , 8 )
Find b by plugging the values of m & the point in
y=mx+b
8 = -2.00 + b
b= 10
m= 1/2
The required equation is , y = 1/ 2 x + 10
multiply by 2
2y=x+20
-x+2y=20
m.ananth@hotmail.ca